OpenStudy (anonymous):

ΔABC is translated 6 units up and 3 units left to create ΔA′B′C′. If vertex A is located at (-1, 2), what are the coordinates of A′ ? (5, -1) (-4, 8) (2, -4) (-7, -1)

undeadknight26 (undeadknight26):

We have (-1, 2) We know it goes up 6 and left 3 So: -1 + 6 = 5 2-3 = -1 (5,-1) is your new point :D

OpenStudy (anonymous):

that was wrong :(

undeadknight26 (undeadknight26):

Oh nvm I see what I did wrong /.- I did it backwards ._.

undeadknight26 (undeadknight26):

-1 = x 2 = y so: -1 - 3 = -4 2 + 6 = 8 (-4,8)

undeadknight26 (undeadknight26):

Sorry...

OpenStudy (anonymous):

(-4,8)